# Identifying compact symmetric objects with high-precision VLBI and Gaia astrometry

### v1

@@ 5 Summary
In this study, we have developed and validated a new method combining Gaia optical astrometry with multi-frequency VLBI observations to identify and characterize CSOs. Our main findings are:
@@
Our innovative Gaia+VLBI approach has decisively classified 32 of 40 CSO candidates, confirming 22 genuine CSOs (17 through direct positional alignment of Gaia with symmetric radio structures, 5 through compelling morphological and spectral evidence despite larger positional offsets), and identifying 10 as core-jet sources. Eight sources remain ambiguous due to significant radio–optical discrepancies. This methodology effectively overcomes the long–standing challenge of detecting faint or absorbed radio cores in young radio galaxies.
@@
We provide robust observational evidence for a two–track evolutionary paradigm in the CSO population. High–power CSOs () demonstrate the potential to evolve into large–scale radio galaxies, exhibiting moderate relativistic hotspot advance speeds (typically ) and systematic spectral gradients. In contrast, lower–power CSOs show characteristics consistent with the Frustrated Radio Galaxy concept, remaining confined within sub-kiloparsec scales with slower expansion speeds () and showing significant environmental impact on their morphology and spectral properties.
@@
The quantitative relationships between radio power, kinematic age, hotspot velocity, and projected source size align remarkably well with theoretical predictions, revealing distinct evolutionary trajectories for high- and low–power CSOs. This confirms our earlier proposed evolutionary model where initial radio power significantly determines a source’s developmental pathway and ultimate fate.
@@
Our comprehensive analysis of CSO properties strongly challenges isolated TDE–driven formation as the primary origin mechanism for most CSOs. The observed extended timescales (up to years), persistent symmetrical jet structures, and substantial energy requirements are inconsistent with typical TDE characteristics but align with standard AGN accretion processes driven by gas infall during galaxy mergers or bar-induced instabilities.
@@
Beyond radio power, we demonstrate that CSO evolution is governed by a complex interplay of environmental factors. The ambient medium density distribution, magneto-ionic conditions in surrounding gas, intermittent accretion episodes, and central engine activity duration collectively shape CSO morphology, spectral signatures, and evolutionary trajectories, explaining the diverse observed characteristics within our sample.
6.
CSOs serve as crucial laboratories for studying AGN feedback mechanisms across different power scales. High–power sources generate significant mechanical feedback on relatively large physical scales, while lower–power sources contribute predominantly to localized heating and ISM disruption in galactic cores, often without developing large–scale radio structures. This diversity reveals multiple channels of AGN feedback relevant to galaxy evolution models.
7.
Future applications of our Gaia+VLBI methodology to larger samples will significantly advance our understanding of radio galaxy evolution. Particularly valuable would be high–sensitivity observations targeting low–power CSOs with next–generation facilities like SKA, which could reveal previously undetected populations and provide deeper insights into AGN feedback, the prevalence of frustrated sources, and complex interplay between radio sources and their host environments throughout cosmic history.
@@ 6 Data availability
The VLBI datasets underlying this article were derived from the public domain in the Astrogeo archive (http://astrogeo.org/).
@@ Acknowledgements.
This work is supported by the National Key R&D Programme of China (2022SKA0120102). TA and WAB acknowledge the support from the Xinjiang Tianchi Talent Program. YKZ is supported by the China Scholarship Council (No. 202104910165), the Shanghai Sailing Program under grant number 22YF1456100, and the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDA0350205). YKZ thanks for the warm hospitality and the helpful comments from Ivy Wong in CSIRO Space&Astronomy in Australia. YKZ thanks for the valuable suggestions and comments from Krisztina Éva Gabányi in Konkoly Observatory, Hungary. SF thanks the Hungarian National Research, Development and Innovation Office (NKFIH, grant no. OTKA K134213) for support. This work was also supported by the NKFIH excellence grant TKP2021-NKTA-64. The authors acknowledge the use of Astrogeo Center database maintained by L. Petrov. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). This work has made use of the NASA Astrophysics Data System Abstract Service, and the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. The National Radio Astronomy Observatory are facilities of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.References
@@
- Bicknell et al. (2003) Bicknell, G. V., Saxton, C. J., Sutherland, R. S., Midgley, S., & Wagner, S. J. 2003, New A Rev., 47, 537
@@
- Charlot (1990) Charlot, P. 1990, AJ, 99, 1309
- Chen et al. (2023a) Chen, S., Laor, A., Behar, E., Baldi, R. D., & Gelfand, J. D. 2023a, MNRAS, 525, 164
- Chen et al. (2023b) Chen, Y.-C., Liu, X., Lazio, J., et al. 2023b, ApJ, 958, 29
@@
- Fomalont (1999) Fomalont, E. B. 1999, in Astronomical Society of the Pacific Conference Series, Vol. 180, Synthesis Imaging in Radio Astronomy II, ed. G. B. Taylor, C. L. Carilli, & R. A. Perley, 301
- Fomalont et al. (2003) Fomalont, E. B., Petrov, L., MacMillan, D. S., Gordon, D., & Ma, C. 2003, AJ, 126, 2562
@@
- Kellermann & Owen (1988) Kellermann, K. I. & Owen, F. N. 1988, in Kellermann, K. I. and Verschuur, G. L., eds, Galactic and Extragalactic Radio Astronomy, 2nd edition, ed. K. I. Kellermann & G. L. Verschuur, Springer-Verlag (Springer-Verlag, Berlin–New York), 563–602
@@
- Kellermann et al. (1994) Kellermann, K. I., Sramek, R. A., Schmidt, M., Green, R. F., & Shaffer, D. B. 1994, AJ, 108, 1163
@@
- Kovalev et al. (2007) Kovalev, Y. Y., Petrov, L., Fomalont, E. B., & Gordon, D. 2007, AJ, 133, 1236
- Kukreti et al. (2023) Kukreti, P., Morganti, R., Tadhunter, C., & Santoro, F. 2023, A&A, 674, A198
@@
- Lindegren et al. (2018) Lindegren, L., Hernández, J., Bombrun, A., et al. 2018, A&A, 616, A2
- Lindegren et al. (2012) Lindegren, L., Lammers, U., Hobbs, D., et al. 2012, A&A, 538, A78
@@
- Petrov et al. (2005) Petrov, L., Kovalev, Y. Y., Fomalont, E., & Gordon, D. 2005, AJ, 129, 1163
- Petrov et al. (2006) Petrov, L., Kovalev, Y. Y., Fomalont, E. B., & Gordon, D. 2006, AJ, 131, 1872
- Petrov et al. (2008) Petrov, L., Kovalev, Y. Y., Fomalont, E. B., & Gordon, D. 2008, AJ, 136, 580
- Petrov et al. (2019) Petrov, L., Kovalev, Y. Y., & Plavin, A. V. 2019, MNRAS, 482, 3023
@@
- Santoro et al. (2020) Santoro, F., Tadhunter, C., Baron, D., Morganti, R., & Holt, J. 2020, A&A, 644, A54
@@
- Shabala & Alexander (2009) Shabala, S. & Alexander, P. 2009, ApJ, 699, 525
- Shen et al. (2021) Shen, Y., Chen, Y.-C., Hwang, H.-C., et al. 2021, Nature Astronomy, 5, 569
- Shepherd (1997) Shepherd, M. C. 1997, Astronomical Society of the Pacific Conference Series, Vol. 125, Difmap: an Interactive Program for Synthesis Imaging, ed. G. Hunt & H. Payne, 77
@@
- Sovers et al. (1998) Sovers, O. J., Fanselow, J. L., & Jacobs, C. S. 1998, Reviews of Modern Physics, 70, 1393
@@
- Tadhunter et al. (2021) Tadhunter, C., Morganti, R., Santoro, F., & Bernhard, E. 2021, Astronomische Nachrichten, 342, 1200
- Taylor et al. (2005) Taylor, G. B., Fassnacht, C. D., Sjouwerman, L. O., et al. 2005, ApJS, 159, 27
@@
- Taylor et al. (1996) Taylor, G. B., Readhead, A. C. S., & Pearson, T. J. 1996, ApJ, 463, 95
@@
- White et al. (2020) White, S. V., Franzen, T. M. O., Riseley, C. J., et al. 2020, PASA, 37, e018
@@
Column (1): source name. Column (2): Equatorial coordinates from Astrogeo (Petrov & Kovalev 2025). Column (3): Spectroscopic or photometric redshifts collected from the NASA/IPAC Extragalactic Database (https://ned.ipac.caltech.edu/) or SIMBAD (https://simbad.u-strasbg.fr/). Column (4) lists the optical identification of the source type from SIMBAD. Q – quasar; G – galaxy. Column (5) lists the references of the sources. Column (6) presents the classification of the radio morphology: CSO – confirmed CSO with Gaia-identified AGN between two radio components; CJ – one-sided core–jet with Gaia-identified AGN at the end of the radio structure; CSO candidate – sources with large optical–radio offset remain as CSO candidates. References: 1 – Peck & Taylor (2000); 2 – An & Baan (2012); 3 – Tremblay et al. (2016).
Name | Comp. | PAvar | Age | ||
(mas yr-1) | (c) | ( yr-1) | (yr) | ||
J0832+1832 | W-E | 0.0060.004 | 0.060.04 | 0.06 | 1200820 |
J0943+1702 | N-S1 | 0.0290.014 | 2.111.00 | 0.050.05 | 15070 |
S2-N | 0.0220.015 | 1.581.08 | 0.070.03 | 240160 | |
J1110+4817 | SW-NE1 | 0.015 | 0.64 | 0.08 | 830 |
J1111+1955 | W-E | 0.2950.062 | 5.551.16 | 0.31 | |
J1158+2450 | NE-C | 0.0010.006 | 0.010.08 | 0.07 | 1160 |
NW-C | 0.0320.007 | 0.420.09 | 0.060.21 | 8020 | |
J1234+4753 | SE-C | 0.0020.003 | 0.060.08 | 0.06 | 1840 |
SE-NW | 0.0040.004 | 0.100.08 | 0.05 | 16801400 | |
J1244+4048 | SW2-C | 0.0250.045 | 1.102.00 | 0.150.25 | 130 |
J1247+6723 | SE-NW | 0.0460.010 | 0.330.07 | 0.030.15 | 16030 |
J1358+4737 | W-E | 0.0210.004 | 0.310.06 | 0.070.10 | 25050 |
J1407+2827 | W-E | 0.0420.003 | 0.220.02 | 0.04 | 16010 |
J1511+0518 | W-E | 0.0100.002 | 0.060.01 | 0.040.09 | 480100 |
J1602+2418 | W-E | 0.0330.006 | 2.540.45 | 0.430.12 | 8010 |
J1815+6127 | SW-NE1 | 0.033 | 1.16 | 0.44 | 160 |
NE2-SW | 0.035 | 1.21 | 0.35 | 200 | |
J1823+7938 | W2-E | 0.0060.008 | 0.090.12 | 0.10 | 1580 |
J1945+7055 | SW2-NE3 | 0.0210.006 | 0.140.04 | 0.02 | 1030270 |
SW1-NE2 | 0.0130.004 | 0.080.03 | 0.06 | 720220 | |
SW2-NE2 | 0.0580.006 | 0.380.04 | 0.03 | 28030 |
Column (1): source name; Column (2): selected component with respect to the reference component used for proper motion measurement; Column (3): radial proper motion; positive values mean the components moving away from each other; Column (4): apparent radial separation speed measured in the units of in the source rest frame; Column (5): position angle change, measured from north to east; Column (6): dynamical age of the component; if a component has negative or negligible proper motion, we used 1- of the separation velocity to derive a lower limit (see An et al. 2012). The fitted proper motion plots are shown in Fig. 7.
Name | Radio power | Projected size | Hotspot velocity | Kinematic age | Spectral index | Ref |

### v2

@@ 5 Summary
In this study, we developed and validated a Gaia+VLBI method to identify and characterize CSOs. Our main findings are listed as follows.
@@
The Gaia+VLBI approach decisively classifies 30 out of 40 candidates: 20 confirmed CSOs (15 via direct Gaia–VLBI alignment; 5 via morphology/spectra despite larger offsets) and ten core–jet sources; ten remain ambiguous. This overcomes the long-standing difficulty of locating faint/absorbed radio cores in young sources. Scaling this Gaia+VLBI framework to larger samples, especially with sensitive SKA-era observations of low-power CSOs, will reveal hidden populations and sharpen constraints on feedback, frustrated sources, and host-galaxy-jet coupling across cosmic time.
@@
The two-track evolutionary paradigm is reinforced by multiple, independent observational correlations presented here. High-power CSOs () show moderate hotspot speeds (typically ), systematic spectral gradients, and potential growth into large radio galaxies. Low-power CSOs remain sub-kiloparsec with slower expansion () and strong environmental imprint, consistent with the frustrated radio galaxy picture. Correlations among radio power, size, hotspot speed, and kinematic age match theoretical expectations, implying distinct trajectories set largely by initial radio power.
@@
The properties of CSOs suggest that isolated TDEs are unlikely to be the dominant origin: long activity, persistent symmetry, and energy budgets in some CSOs instead point to standard AGN fueling.
@@
Beyond power, evolution reflects environment and duty cycle: ambient density and magneto-ionic conditions, intermittent accretion, engine lifetime shape morphology, spectra, and growth, hence the observed diversity of observational characteristics.
@@
Compact symmetric objects are laboratories for AGN feedback: high-power systems deliver mechanical work on larger scales; low-power systems heat and disturb the central ISM without developing large lobes.
@@ 6 Data availability
The VLBI datasets underlying this article were derived from the public domain in the Astrogeo archive (http://astrogeo.org/). The materials originally presented as appendices are available as online supplementary material at Zenodo (DOI: https://doi.org/10.5281/zenodo.17278723). Figure and table numbering in the supplement follows the S-scheme (e.g. Fig. S1, Table S1).
@@ Acknowledgements.
We thank the referee for the constructive comments that improved the paper. TA and WAB are grateful to Anthony C. Readhead for his constructive discussions and valuable suggestions. This work is supported by the National Key R&D Program of China (2022SKA0120102). This work used resources of China SKA Regional Centre prototype funded by the Ministry of Science and Technology of the People’s Republic of China and the Chinese Academy of Sciences (An et al., 2019, 2022). TA and WAB acknowledge the support from the Xinjiang Tianchi Talent Program. YKZ is supported by the China Scholarship Council (No. 202104910165), the Shanghai Sailing Program under grant number 22YF1456100, and the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDA0350205). YKZ thanks for the warm hospitality and the helpful comments from Ivy Wong in CSIRO Space&Astronomy in Australia. YKZ thanks for the valuable suggestions and comments from Krisztina Éva Gabányi in Konkoly Observatory, Hungary. SF thanks the Hungarian National Research, Development and Innovation Office (NKFIH, grant no. OTKA K134213) for support. This work was also supported by the NKFIH excellence grant TKP2021-NKTA-64. The authors acknowledge the use of Astrogeo Center database maintained by L. Petrov. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). This work has made use of the NASA Astrophysics Data System Abstract Service, and the NASA/IPAC Extragalactic Database (NED), which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.References
@@
- An et al. (2022) An, T., Wu, X., Lao, B., et al. 2022, Science China Physics, Mechanics, and Astronomy, 65, 129501
- An et al. (2019) An, T., Wu, X.-P., & Hong, X. 2019, Nature Astronomy, 3, 1030
@@
- Chen et al. (2023) Chen, S., Laor, A., Behar, E., Baldi, R. D., & Gelfand, J. D. 2023, MNRAS, 525, 164
@@
- DESI Collaboration et al. (2025) DESI Collaboration, Abdul-Karim, M., Adame, A. G., et al. 2025, arXiv e-prints, arXiv:2503.14745
@@
- Kosogorov et al. (2024) Kosogorov, N. A., Kovalev, Y. Y., Perucho, M., & Kovalev, Y. A. 2024, MNRAS, 528, 1697
@@
- Taylor et al. (1996a) Taylor, G. B., Readhead, A. C. S., & Pearson, T. J. 1996a, ApJ, 463, 95
- Taylor et al. (1996b) Taylor, G. B., Vermeulen, R. C., Readhead, A. C. S., et al. 1996b, ApJS, 107, 37
@@
- White et al. (2020a) White, S. V., Franzen, T. M. O., Riseley, C. J., et al. 2020a, PASA, 37, e018
- White et al. (2020b) White, S. V., Franzen, T. M. O., Riseley, C. J., et al. 2020b, PASA, 37, e017
@@ Appendix A Supplementary tables
| Name | Radio power | Projected size | H–H velocity | Kinematic age | Spectral index | Ref |